What is the price of a share of stock if the beta is 2, its next dividend is projected to be $ 2 and its growth rate is expected to be a constant 9 %, assuming the market return is 16% and the risk-free rate is 7 %?
Answer #1

required return= risk-free rate +Beta*(MArket rate- risk-free rate )

=7+2*(16-7)

=25%

Current price=D1/(Required return-Growth rate)

=2/(0.25-0.09)

which is equal to

=$12.5